OpenStudy (brydenwright):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (mathmate):

hint: \(\Large \cos(\theta)=\frac{A\circ B}{||A|| \times ||B||}\) i.e. dot product of two vectors divided by the product of the magnitudes.